Grade 12th pass8 grade maths

For each of the following numbers, find the smallest whole number by which it should be multiplied so as to get a perfect square number. Also find the square root of the square number so obtained. i. 252

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

252 = 2×2×3×3×7
= (2×2)×(3×3)×7
Here, 7 cannot be paired.
∴ We will multiply 252 by 7 to get perfect square.Newnumber=252×7=1764


1764 = 2×2×3×3×7×7
⇒ 1764 = (2×2)×(3×3)×(7×7) ⇒ 1764 =22×32×72
⇒ 1764 = (2×3×7)2⇒√1764=2×3×7=42



Thanks